Select the correct answers
 
A particle is projected vertically upwards in vacuum with a speed u. Then
 
A) When it rises to half its maximum height, its speed becomes u/2.
B)When it rises to half its maximum height, its speed becomes u√2
C)The time taken to rise to half its maximum height is half the time taken to reach its maximum height.
D)the time taken to rise to three-fourth of its maximum height is half the time taken to reach its Maximum height

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Maximum height=\dfrac{u_y^2}{2g}
Time taken to reach maximum height=t_0=\dfrac{u_y}{g}

For motion from groudn to 3/4th the maximum height,

\dfrac{3}{8}\dfrac{u_y^2}{g}=u_yt-\dfrac{1}{2}gt^2
\implies t=\dfrac{u_y}{2g}=\dfrac{t_0}{2}